what would be the principal ballon payment (that is excluding the fixed monthly payment) at the end of 5 years on a $150,000 loan with monthly payments based on a 30-year term at 12 percent compounded monthly? the answer is 146,495 but i dont know how to get this answer.
Answer #1

Monthly payment of loan pv =150000 r= 12%/12 =1% n =12*30 =360

PMT = Pv*r / 1-(1.01)^-360

150000*1% / 1-(1.01)^-360 = 1542.92

Baloon payment = pv*(1+r)^n - pmt*[1-(1+r)^n -1 / r]

150000*(1.01)^60 - 1542.92*[(1.01)^60 -1 / .01 ]

272504.505 - 126009.767 = 146494.738 = 146495